OpenStudy (anonymous):

A gas company has the following rate schedule for natural gas usage in single-family residences: Monthly service charge $8.80 Per therm service charge 1st 25 therms $0.6686/therm Over 25 therms $0.85870/therm What is the charge for using 25 therms in one month? What is the charge for using 45 therms in one month? Construct a function that gives the monthly charge C for x therms of gas.

OpenStudy (anonymous):

ganeshia can you help please

ganeshie8 (ganeshie8):

monthly service charge is flat. you have to pay it no matter how many therms you use.

ganeshie8 (ganeshie8):

What is the charge for using 25 therms in one month? monthly service charge + something

ganeshie8 (ganeshie8):

since you've used only 25 therms, they'll charge you $0.6686/therm

ganeshie8 (ganeshie8):

so, total charge for using 25 therms in one month = 8.8 + 25*0.6686 = ?

OpenStudy (anonymous):

16.715+8.8 = 25.52

ganeshie8 (ganeshie8):

correct, next q

ganeshie8 (ganeshie8):

What is the charge for using 45 therms in one month? 45 = 25 + 20 so for first 25 therms they'll charge u .6686/therm, and rest of the 20 they'll charge u 0.85870/therm

ganeshie8 (ganeshie8):

8.8 + 25*.6686 + 20*.85870

OpenStudy (anonymous):

25.52+17.174 = 42.694

ganeshie8 (ganeshie8):

thats right ! next is the tricky part

OpenStudy (anonymous):

for this i have written this C(x) = { 8.8 + 0.6686x if 0 ≤ x ≤ 25 { 25.515 + 0.133219(x - 25) if x > 25

ganeshie8 (ganeshie8):

how did u get 0.133219 ?

ganeshie8 (ganeshie8):

it should be, 0.85870 right ?

OpenStudy (anonymous):

i thought of subtracting it from 0.6686

ganeshie8 (ganeshie8):

i see you did most of it real good,

OpenStudy (anonymous):

can you help me with some similar questions Please

ganeshie8 (ganeshie8):

C(x) = { 8.8 + 0.6686x if 0 ≤ x ≤ 25 { 25.515 + 0.85870(x - 25) if x > 25

ganeshie8 (ganeshie8):

we dont have to subtract cuz, you accounted for the first 25 therms in 25.515 : 25.515 = 8.8 + 25*.6686

ganeshie8 (ganeshie8):

so rest of the therms, we need to multiply flat with .85870

ganeshie8 (ganeshie8):

sure post them il be happy to help if i can :D

OpenStudy (anonymous):

http://prntscr.com/1504bo

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

v = 12 falls in 2nd ? so use it plugin v=12, t = 15

OpenStudy (anonymous):

ganiesha we have to answer it in 3 parts

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

where are the 3 parts

ganeshie8 (ganeshie8):

the function has 3 parts, but only one we can use. so solution doesnt require 3 parts